# Lexical categories of stem-forming roots in Mapudüngun verb forms

### v2

@@ 4.2 Tables
Table 3: Verbal Roots Reclassified As Other PoS
Root Extracted
Suffixes
Initial
Cat.
Final
Cat.
Definition
ad- VI N the visible or characteristic aspects of something or someone, or of a group
af- VI N end
afü- VI Aj cooked, ripe
allush- VI Aj lukewarm
alü- VI Av very, much
angim- VT N jerky
angka- VI N middle, half
antü- VI N Sun, day, time
are- VI Aj hot, warm
ariñ- VI Aj burnt, smoked
arof- VI N sweat
asul- VI Aj blue
awi- VI Aj hot
awka- VI Aj rebellious, wild
aylen- VI N ember(s)
aywiñ- VI N shade
chadi- VI N salt
chafo- VI N cough
chali- VI N greeting
challwa- VI N fish
chape- VT N braid (in the hair)
che- VI N person, people
chedkuy- VT N father/son in law (of a man)
chilla- VT N saddle, chair
chod- VI Aj yellow
chum- VT Int how, what
chüngküd- VI Aj circular, rounded
dakel- VT N commission, agreement
dew- VI Av already, since, because, no more, only, finally
domo- VI N woman
[…]
duam- VT N matter, issue, need, intention
dumdum- VI N penumbra
düngu- VT N issue, matter, subject, concern, word, language
[…]
elfal- VT N commission
eñum- VI Aj comfortingly/pleasantly warm
epew- VI N tale, story, history, narration
fa- VI Dem this, here
fe- VI Dem that, there
fente- VT Av so, much
fentre- VI Av very, much
fey- VI Dem this, that, she, he
filla- VI N scarcity, lack
fitrun- VI N smoke
fücha- VI Aj big, old
fülang- VI Aj white
funa- VI Aj rotten
füre- VI Aj spicy, hot
futrül- VT N heap, pile
füw- VT N thread, wool
ina- VI Av close
ingka- VT N defender
kachu- VI N grass, herb
kafe- VI N coffee
kallfü- VI Aj blue, purple, violet
kangka- VT Aj roasted
karü- VI Aj green
kashü- VI Aj gray
katrü- VT Aj cut off, interrupted
kawe- VI N oar
kellu- VT N help
kelü- VI Aj red
kim- VT Aj wise, knower
kochü- VI Aj sweet
kofke- VT N bread
kolü- VI Aj rust, reddish-brown (colour)
kompañ- VT N partner
koñoll- VI Aj purple, violet
koñü- VI N baby, newborn
korü- VI N soup, broth
kotrü- VI Aj salty, sour, bitter
kowkow- VI N owl
küdaw- VI N work
kude- VI N bet
[…]
küllay- -tu- VT N quillaya, quillay soap
külü- VI Aj tilted, slanted
küme- VI Aj good
kunaw- VI Aj inflated, floating
kupül- VT Aj fixed, secure
kuram-b VI N egg
küri- VI Aj lustful, lecherous
kurü- VI Aj black
kürüf- VI N wind
kushe- VI Aj elderly, old (woman)
kutran- VI N illness, disease, pain, suffering
kütrüng- VT N bundle, bunch
küyen- VI N moon, month
la- VI Aj dead, deceased
lantra- VI Aj bulging, thick
lawen- VT N medicine (medicinal herb)
lef- VI Aj fast (going/moving)
leg- VI Aj certain, accurate, correct
lichi- VI N milk
lif- VI Aj clean
[…]
llag- VT Aj half
llekü- VI Av close
llitu- VI Aj initial, first
[…]
llüngüd- VT Aj flat, level
llüpañ- VT Aj lying, reclining
lüf- VI Aj scorched, roasted
lüp-d -üm- VT Aj scorched, roasted
lüykü- VI N drop
machi- VI N shaman
[…]
mafü- VT N dowry
[…]
mangkü- VT N kick, boot
mañum- VT N gratitude, thankfulness
mawün- VT N rain
may- VT Conj yes
me- VI N excrement, scum, dross
melkay- VI Aj slippery
[…]
minggako- VI N communal/community work
misha- VT N dining companion
[…]
motri- VI Aj fat
moy- VI N boil
moyo- VT N breast
muday- VI N chicha
[…]
müna- -le- VI Av many, much
mupi- VI Aj true, genuine
müpü- VI N wing
[…]
mütrüm- VT N call
nag- VI Av down, below, under, beneath
[…]
ngeñi- -ka- VI Aj urgent, pressing
ngoyma- VT Aj unproductive, neglected
ngüdi- VT N obstruction
ngülam- VT N advice, tip
ngüñü- VI N hunger
nor- VI Aj straight (in a physical or figurative sense), correct
notu- -ka- VT N obstinacy, blindness
nüla- VT Aj open
nümü- VI N smell, odour
[…]
nürüf- VT Aj closed
nütram- VT N recount, conversation, story
nüyün- VI N tremor, quake
ñochi- VI Aj peaceful, calm
ñum- VI N upper millstone
paf- VI N boil, abscess
[…]
pali- VI N paline ball
panü- VT N fathom
papel- VT N paper
[…]
perimo- -ntu- VI N omen
perkü- VI N fart
pewma- VI N daydream, dream
pichi- $similar-to\sim püchü- simi$$similar-to\sim püt$ü- VI Aj little
pidku- VT Aj stew (referring to a dish typically involving legumes cooked in a pot)
piku- VI N North
pilu- VI Aj deaf
pire- VI N snow, hail
[…]
pitru- VI N scabies
pod- VI Aj dirty
pofre- VI Aj poor
poy- VI N tumour, swelling
püd- VI Aj thick, dense, condensed
pülle- VI Av close, next
püntü- VI N piece, chunk
püñeñ- VI N daughter or son (to the mother or her sister)
püro- VT N knot
pütre- VT Aj blazing, scorching, burning, hot
pütrü- VI Av in abundance, a lot, plentifully
raküm- VT N closure, wall, protection
rangiñ- VI N middle, half
re- -le- VI Aj alone, exclusive, pure
riku- VI Aj rich (it may be related to rükü- (further down)
ruka- VT N house
rükü- VI Aj mean, petty, greedy, miserly, tight-fisted, stingy
rümpel- VI Aj jealous
[…]
rutra- -tu- VT N pinch
tonon- VT N weft
traf- VI Aj together, gathered, joined, fitted, interlocked
traf-iya-f VI Av last night, since last night’s fall
tralka- VI N thunder, shotgun
tran- VI Aj fallen (especially referring to trees)
trangliñ- VI N frost
trapel- VT N cord, string
trar- VI N pus
[…]
traytray- VI N gargle, or similar gargling sound
trem- VI Aj grown, adult, mature
tremo- VT Aj healthy, robust, splendid
trintri- VI Aj curly
tripantu- VI N year
tromü- VI N cloud
tror- VI N foam
trüfon- VI N cough
trufür- VI N dust
[…]
trülke- VT N leather
[…]
trutruka- VI N a wind instrument like a long cornet
tute- VT Aj correct, accurate
üküm- VI N silence
ül- VI N song, singing
[…]
üna- VI N itch
üng- -üm- VT AV while
[…]
üyag- -tu- VT Aj both
wachi- VT N trap
waw- VI N valley, lowland, low-lying area
wechod- VI N hole
wedwed- VI Aj mischievous
welu- VI Aj alternative, alternating
wentru- VI N man
wew- VT N victory, triumph, gain, profit
[…]
wif- VI Aj wide, broad
[…]
wilüf- VI Aj shiny
wim- VI Aj tame, accustomed
wima- VT N rod, staff
witral- VT N warp, loom
wün- VI N mouth, dawn, sunrise
wüne- VI Aj fore, anterior, superior
wütre- VI Aj cold
yafü- VI Aj hard, firm
yall- VT N son
[…]
yochi- VI Aj tight, fitted
yung- -üm- VT Aj pointed, sharp
a. Küllay is a tree endemic to the central region of Chile; scientific name Quillaja saponaria, also known as ‘soapbark tree’.
b. In conjugated (verbalised) forms, the root kuram ‘egg’ is alternatively analysed as -NN.kürastone+CA.m34… by Düngupeyüm, that is to say, as kura-m, a form that may be interpreted as ‘caused to be a stone’, which may well be the origin of kuram ‘egg’.
[…]
d. The causative suffix -(ü)m- triggers a radical change in some roots, causing those ending in f to shift to p (see Chandía, 2025, E.06), indicating that this root (lüp-) is the same as the one above (lüf-).
e. Palin is a sport similar to field hockey, referred to as chueca by Spaniards and Creoles.
[…]
Table 4: Roots That Causativise With -(ü)m-
@@
1. ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
2. afü Aj cooked, ripe afü-m- to cook K. S. G.
3. ali VI to get heated up ali-m- to heat up S.
4. angkü VI to get dry (up), to be sterile, to be arid angkü-m- to dry (up) K. S.
5. anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
6. are Aj hot, warm are-m- to warm up S. G.
7. chaw N father, man chaw-üm-b to incubate K. S.
8. che N person, people, human being che-m- to relatec K.
9. chong VI to go out, to extinguish, to dry up chong-üm- to turn off, to extinguish, to dry S.
@@
11. fa Dem this, here fa-m- to be like this K. S.
12. fay VI to ferment fay-üm- to [make] ferment S.
13. fe Dem that, there fe-m- to be like that K. S.
14. fey VI to get adapted, to fit fey-üm- to adapt, to fit, to insert S.
15. firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
16. fül Av close fül-üm- to bring closer, to approximate, to bring near K. S.
17. füy VI to squeeze (itself), to hold on (to itself) füy-üm- to tighten, to hold K.
18. illku VI to get angry illku-m- to contradict, to make angry K.
19. kofi VI to get heat up, to get overheat kofi-m- to heat, to overheat S.
20. kon VI to enter, to begin kon-üm- to make come in, to add K. G.
21. kudu VI to lie (down), to go to bed kudu-m- to lay out K.
22. küme Aj good, nice, pleasant küme-m- to add value, to improve K.
@@
23. kümpu N piece, chunk, bit kümpu-m- to chop, to cut into pieces, to tear apart, to shred K.
24. küpa VI to come küpa-m- to bring, to introduce K.
25. la Aj dead, deceased lang-üm- to kill K. S. G.
26. laf Aj flat, level lap-üm- to stretch, to extend S.
27. law VI to lose hair/feathers law-üm- to shave, to shear, to pluck S.
@@
28. lef Aj fast (on going or moving) lep-üm- to make run (animals), to make flee, to make turn away S. G.
29. (l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
30. lif Aj clean lip-üm- to clean K.
31. llangkü VI to fall (down) llangkü-m- to drop, to lose K. S.
@@
32. lleg VI to grow (up), to come up llek-üm- to raise, to cultivate, to grow (plants), to bring out S. G.
[…]
34. llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
35. lüf Aj burned, roasted p-üm- to burn, to set fire, to roast S. G.
36. lüg Aj white lüg-üm- to whiten, to bleach, to lighten (in color) S.
37. lulu N noise, crash, upheaval lulu-m- to make noise, to cause crashing K.
38. nag Av down, underneath nak-üm- to go down, to lower, to descend K. S. G.
[…]
40. neng VI to move (oneself) neng-üm- to move, to shake, to shudder, to direct K.
41. neyküe VI to unleash, to get free neykü-m- to release, to let go, to free S.
@@
42. ngül VI to get gathered, to come together, to join ngül-üm- to gather, to bring together, to collect, to harvest K. S.
@@
43. nong VI to get broken down, to get damaged, to get scorched nong-üm- to spoil, to damage, to scorch K.
44. nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
45. ñam VI to get lost, to get astray, to get loose ñam-üm- to lose, to release, to let go, to make disappear S. G.
46. ñif VI to dry out ñip-üm- to dry S.
47. ñom Aj meek, gentle, tame, docile, obedient ñom-üm- to tame, to domesticate, to tame K. S.
48. pelof N light, clarity pelo-m- to illuminate, to light up, to turn on the light S.
@@
49. pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
50. piwü VI to dry out piwü-m- to dry (up) K. S.
51. pod N dirt, filth, uncleanliness pod-üm- to dirty, to soil S.
52. poñü N potato poñü-m- feed with potatoes K.
@@
53. püd Aj thick, dense, scattered, dispersed püd-üm- to spread, to scatter, to thicken, to extend, to spread out, to stretch, to ration, to distribute K. S.
54. pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
55. püna VI to get stuck, to get glued, to get adhered püna-m- to stick, to glue, to adhere S.
@@
56. püra VI to ascend, to go up, to climb püra-m-g to lift, to raise, to pick up, to keep in mind, to remember, to make evident, to highlight K. S. G.
57. püsha VI to run out, to be finished püsha-m- to finish, to end, to spend, to use up, to exhaust G.
58. pütre Aj burning, red-hot pütre-m- to smoke, to fumigateh K.
@@
59. puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
60. rüngü VI to be pulverised, to be ground, to be crushed rüngü-m- to grind, to mill, to crush S.
@@
61. traf Aj together, joined, gathered, fitted trap-üm- to gather, to collect, to put together, to grab, to catch, to pick up K. S.
@@
62. tral VI to get hit, to get bumped, to get dented, to get bruised, to collide, to crash, to clash tral-üm- to hit, to strike, to beat K.
@@
63. tran Aj fallen (especially referring to trees) tran-üm- to make fall, to knock down, to knock over, to topple, to lay down G.
64. trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
65. trem Aj grown, adult, mature trem-üm- to raise, to breed K. S. G.
66. trof VI to burst, to break, to crack, to snap trop-üm- to split, to break S.
67. trong Aj thick, dense trong-üm- to roof (houses)i K.
@@
68. trür Aj even, level, tight, composed trür-üm- to equalise, to compensate, to balance, to indemnify K. S.
69. tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
70. üng Av while üng-üm- to wait K.
[…]
72. ürkü VI to get tired ürkü-m- to tire, to tire out G.
[…]
74. üy N name üy-üm- to name, to denominate S.
75. wadkü VI to boil wadkü-m- to boil, to bring to a boil K. S. G.
76. wallü VI to get surrounded, to get encompassed, to get completed wallü-m- to tell or recount in detail K.
77. wef VI to appear, to show itself wep-üm- to make appear, to find, to show G.
78. welli VI to get empty welli-m- to empty, to vacate K. S.
79. wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
80. wüda VI to divide, to split up wüda-m- to divide, to distribute, to share, to split K. S.
81. wünü VI to get stretched, to get elongated wünü-m- to stretch, to lengthen S.
82. yall N son (of a man) yall-üm- to multiply, to engender S.
83. yewe VI to feel embarrassed, to be ashamed yewe-m- to embarrass G.
84. yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
a. In the ‘‘Collected from’’ column, the reference work containing examples with the corresponding root is identified. K. stands for Paskwal Koña in Wilhelm et al., 1930; S. for Smeets; and G. for Golluscio, 2007.
@@
b. The Mapuche use the term chaw-üm- ‘fathering’, or more literally ‘to cause to father’, to refer to ‘incubating eggs’.

### v3

@@ 4.2   Tables
?tablename? 3: Verbal Roots Reclassified as Other Parts of Speech
Root Extracted Initial Final Definition
Suffixes Cat. Cat.
ad- IV N the visible or characteristic aspects of something or someone, or of a group
af- IV N end
afü- IV Aj cooked, ripe
allush- IV Aj lukewarm
alü- IV Av very, much
angim- TV N jerky
angka- IV N middle, half
antü- IV N Sun, day, time
are- IV Aj hot, warm
ariñ- IV Aj burnt, smoked
arof- IV N sweat
asul- IV Aj blue
awi- IV Aj hot
awka- IV Aj rebellious, wild
aylen- IV N ember(s)
aywiñ- IV N shade
chadi- IV N salt
chafo- IV N cough
chali- IV N greeting
challwa- IV N fish
chape- TV N braid (in the hair)
che- IV N person, people
chedkuy- TV N father/son in law (of a man)
chilla- TV N saddle, chair
chod- IV Aj yellow
chum- TV Int how, what
chüngküd- IV Aj circular, rounded
dakel- TV N commission, agreement
dew- IV Av already, since, because, no more, only, finally
domo- IV N woman
doy- $\sim$ yod- TV Av more
duam- TV N matter, issue, need, intention
dumdum- IV N penumbra
düngu- TV N issue, matter, subject, concern, word, language
echid- $\sim$ echiw- IV N sneeze
elfal- TV N commission
eñum- IV Aj comfortingly/pleasantly warm
epew- IV N tale, story, history, narration
fa- IV Dem this, here
fe- IV Dem that, there
fente- TV Av so, much
fentre- IV Av very, much
fey- IV Dem this, that, she, he
filla- IV N scarcity, lack
fitrun- IV N smoke
fücha- IV Aj big, old
fülang- IV Aj white
funa- IV Aj rotten
füre- IV Aj spicy, hot
futrül- TV N heap, pile
füw- TV N thread, wool
ina- IV Av close
ingka- TV N defender
kachu- IV N grass, herb
kafe- IV N coffee
kallfü- IV Aj blue, purple, violet
kangka- TV Aj roasted
karü- IV Aj green
kashü- IV Aj gray
katrü- TV Aj cut off, interrupted
kawe- IV N oar
kellu- TV N help
kelü- IV Aj red
kim- TV Aj wise, knower
kochü- IV Aj sweet
kofke- TV N bread
kolü- IV Aj rust, reddish-brown (colour)
kompañ- TV N partner
koñoll- IV Aj purple, violet
koñü- IV N baby, newborn
korü- IV N soup, broth
kotrü- IV Aj salty, sour, bitter
kowkow- IV N owl
küdaw- IV N work
kude- IV N bet
kufü- $\sim$ kufüll- VI/TV Aj spoiled, damaged (due to the effect of heat)
küllay- -tu TV N quillaya, quillay soap
külü- IV Aj tilted, slanted
küme- IV Aj good
kunaw- IV Aj inflated, floating
kupül- TV Aj fixed, secure
kuram-b IV N egg
küri- IV Aj lustful, lecherous
kurü- IV Aj black
kürüf- IV N wind
kushe- IV Aj elderly, old (woman)
kutran- IV N illness, disease, pain, suffering
kütrüng- TV N bundle, bunch
küyen- IV N moon, month
la- IV Aj dead, deceased
lantra- IV Aj bulging, thick
lawen- TV N medicine (medicinal herb)
lef- IV Aj fast (going/moving)
leg- IV Aj certain, accurate, correct
lichi- IV N milk
lif- IV Aj clean
ling-ar $\sim$ lüng-af $\sim$ lüg- $\sim$ lig- -ar, -afc IV Aj white, clear
llag- TV Aj half
llekü- IV Av close
llitu- IV Aj initial, first
lliwa- $\sim$ lluwa- -tu TV N soothsayer, seer
llüngüd- TV Aj flat, level
llüpañ- TV Aj lying, reclining
lüf- IV Aj scorched, roasted
lüp-d -üm TV Aj scorched, roasted
lüykü- IV N drop
machi- IV N shaman
madom- $\sim$ masom- tuku- TV N ember, charred log
mafü- TV N dowry
mallma- $\sim$ malma- IV Aj boastful, arrogant, pretentious
mangkü- TV N kick, boot
mañum- TV N gratitude, thankfulness
mawün- TV N rain
may- TV Conj yes
me- IV N excrement, scum, dross
melkay- IV Aj slippery
meñku- $\sim$ menku- TV N load, burden
minggako- IV N communal/community work
misha- TV N dining companion
mollfü- $\sim$ mollfüñ- IV N blood
motri- IV Aj fat
moy- IV N boil
moyo- TV N breast
muday- IV N chicha
mukur- $\sim$ mukud- IV Aj sour, bitter
müna- -le IV Av many, much
mupi- IV Aj true, genuine
müpü- IV N wing
mürke- $\sim$ murke- TV N toasted flour
mütrüm- TV N call
nag- IV Av down, below, under, beneath
nel- $\sim$ ney- IV Aj loose, unfastened
ngeñi- -ka IV Aj urgent, pressing
ngoyma- TV Aj unproductive, neglected
ngüdi- TV N obstruction
ngülam- TV N advice, tip
ngüñü- IV N hunger
nor- IV Aj straight (in a physical or figurative sense), correct
notu- -ka TV N obstinacy, blindness
nüla- TV Aj open
nümü- IV N smell, odour
nüngay- $\sim$ ningay- IV N anger, fury, rage, impatience
nürüf- TV Aj closed
nütram- TV N recount, conversation, story
nüyün- IV N tremor, quake
ñochi- IV Aj peaceful, calm
ñum- IV N upper millstone
paf- IV N boil, abscess
pakar- $\sim$ pakaka- IV N a large toad
pali- IV N paline ball
panü- TV N fathom
papel- TV N paper
payun- $\sim$ payuin- $\sim$ payum- $\sim$ payuñ- $\sim$ payun’ IV N beard
perimo- -ntu IV N omen
perkü- IV N fart
pewma- IV N daydream, dream
pichi- $\sim$ püchi- $\sim$ püchü- $\sim$ püti- $\sim$ pütü- IV Aj little
pidku- TV Aj stew (referring to a dish typically involving legumes cooked in a pot)
piku- IV N North
pilu- IV Aj deaf
pire- IV N snow, hail
pitra- $\sim$ pitraw- IV N blister, callus
pitru- IV N scabies
pod- IV Aj dirty
pofre- IV Aj poor
poy- IV N tumour, swelling
püd- IV Aj thick, dense, condensed
pülle- IV Av close, next
püntü- IV N piece, chunk
püñeñ- IV N daughter or son (to the mother or her sister)
püro- TV N knot
pütre- TV Aj blazing, scorching, burning, hot
pütrü- IV Av in abundance, a lot, plentifully
raküm- TV N closure, wall, protection
rangiñ- IV N middle, half
re- -le IV Aj alone, exclusive, pure
riku- IV Aj rich (it may be related to rükü- further down)
ruka- TV N house
rükü- IV Aj mean, petty, greedy, miserly, tight-fisted, stingy
rümpel- IV Aj jealous
runa- $\sim$ truna- TV N two-handed handful
rutra- -tu TV N pinch
tonon- TV N weft
traf- IV Aj together, gathered, joined, fitted, interlocked
traf-iya-f IV Av last night, since last night’s fall
tralka- IV N thunder, shotgun
tran- IV Aj fallen (especially referring to trees)
trangliñ- IV N frost
trapel- TV N cord, string
trar- IV N pus
trari- $\sim$ trarü- TV N ribbon, tape, belt
traytray- IV N gargle, or similar gargling sound
trem- IV Aj grown, adult, mature
tremo- TV Aj healthy, robust, splendid
trintri- IV Aj curly
tripantu- IV N year
tromü- IV N cloud
tror- IV N foam
trüfon- IV N cough
trufür- IV N dust
trüg- $\sim$ trow- $\sim$ trig- IV N crack
trülke- TV N leather
trümfül- $\sim$ trünfül- $\sim$ trüfül- IV Aj twisted, bent
trür- $\sim$ mür- $\sim$ ür- IV N pair, couple
trutruka- IV N a wind instrument like a long cornet
tute- TV Aj correct, accurate
üküm- IV N silence
ül- IV N song, singing
üllüf- $\sim$ illüf- IV Aj unfortunate, miserable, eventful
uma- $\sim$ umañ- $\sim$ umag- $\sim$ umaw- IV N sleepiness, drowsiness
üna- IV N itch
üng- -üm TV Av while
ütrüf- $\sim$ utruf- $\sim$ ütüf- $\sim$ itrüf- TV N throw (in a game)
üy- $\sim$ güy- IV N name
üyag- -tu TV Aj both
wachi- TV N trap
waw- IV N valley, lowland, low-lying area
wechod- IV N hole
wedwed- IV Aj mischievous
welu- IV Aj alternative, alternating
wentru- IV N man
wew- TV N victory, triumph, gain, profit
wichü- $\sim$ wüchü- $\sim$ wüchür- TV Aj twisted, crooked
widi- $\sim$ widü- $\sim$ wüdü- TV N clay vessel/container, pottery
wif- IV Aj wide, broad
wikef- $\sim$ wikür- $\sim$ wünfü- $\sim$ wingür- IV/TV Aj torn, broken
wilüf- IV Aj shiny
wim- IV Aj tame, accustomed
wima- TV N rod, staff
witral- TV N warp, loom
wün- IV N mouth, dawn, sunrise
wüne- IV Aj fore, anterior, superior
wütre- IV Aj cold
yafü- IV Aj hard, firm
yall- TV N son
yiwül- $\sim$ yiwüll- TV N shuttle (referring to the loom)
yochi- IV Aj tight, fitted
yung- -üm TV Aj pointed, sharp
a Küllay is a tree endemic to the central region of Chile; scientific name Quillaja saponaria, also known as ‘soapbark tree’.
b In conjugated (verbalised) forms, the root kuram ‘egg’ is alternatively analysed as -nn.küra_stone+ca.m34… by Düngupeyüm, that is to say, as kura-m, a form that may be interpreted as ‘caused to be a stone’, which may well be the origin of kuram ‘egg’.
c In this case (ling-ar-, lüng-af-), the suffixes were not extracted but are simply presented here. They appear to be suffixes (-ar-, -af-) that imply a semantic change (direction) when added to the base form. They might be unproductive suffixes that have become lexicalised with these roots, or they could be mere phonological fillers. No references to these elements were found in the various studies on Mapudüngun. However, it should be noted that a radical change occurs in the root lüg- $\sim$ lig-, similar in nature to that explained in the companion valency article (Chandía, 2026, E.06), though this is triggered by a suffix about which we have no information.
d The causative suffix -(ü)m triggers a radical change in some roots, causing those ending in f to shift to p (see Chandía, 2026, E.06), indicating that this root (lüp-) is the same as lüf- above.
e Palin is a sport similar to field hockey, referred to as chueca by Spaniards and Creoles.
f trafiya $\sim$ trafüya is a lexicalised compound form, likely derived from traf- ‘to come together’ and wiya $\sim$ wüya ‘yesterday’, meaning ‘the coming together of yesterday [with today]’, referring to the period from the previous evening to the following morning, ‘last night’.
?tablename? 4: Roots That Causativise With -(ü)m
@@
1. ad N form, place, familiar, face, image, custom ad-üm to be experienced/skilled (in something) K. S.
2. afü Aj cooked, ripe afü-m to cook K. S. G.
3. ali IV to get heated up ali-m to heat up S.
4. angkü IV to get dry (up), to be sterile, to be arid angkü-m to dry (up) K. S.
5. anü IV to get settled, to get sit anü-m to settle, to set, to put, to place K. S. G.
6. are Aj hot, warm are-m to warm up S. G.
7. chaw N father, man chaw-ümb to incubate K. S.
8. che N person, people, human being che-m to relatec K.
9. chong IV to go out, to extinguish, to dry up chong-üm to turn off, to extinguish, to dry S.
@@
11. fa Dem this, here fa-m to be like this K. S.
12. fay IV to ferment fay-üm to [make] ferment S.
13. fe Dem that, there fe-m to be like that K. S.
14. fey IV to get adapted, to fit fey-üm to adapt, to fit, to insert S.
15. firkü Aj fresh, cool firkü-m to cool, to warm slightly, to take the chill off K.
16. fül Av close fül-üm to bring closer, to approximate, to bring near K. S.
17. füy IV to squeeze (itself), to hold on (to itself) füy-üm to tighten, to hold K.
18. illku IV to get angry illku-m to contradict, to make angry K.
19. kofi IV to get heat up, to get overheat kofi-m to heat, to overheat S.
20. kon IV to enter, to begin kon-üm to make come in, to add K. G.
21. kudu IV to lie (down), to go to bed kudu-m to lay out K.
22. küme Aj good, nice, pleasant küme-m to add value, to improve K.
@@
23. kümpu N piece, chunk, bit kümpu-m to chop, to cut into pieces, to tear apart, to shred K.
24. küpa IV to come küpa-m to bring, to introduce K.
25. la Aj dead, deceased lang-üm to kill K. S. G.
26. laf Aj flat, level lap-üm to stretch, to extend S.
27. law IV to lose hair/feathers law-üm to shave, to shear, to pluck S.
@@
28. lef Aj fast (on going or moving) lep-üm to make run (animals), to make flee, to make turn away S. G.
29. (l)el IV to let oneself, to become loose, to detach (l)el-üm to release, to let go, to kill K.
30. lif Aj clean lip-üm to clean K.
31. llangkü IV to fall (down) llangkü-m to drop, to lose K. S.
@@
32. lleg IV to grow (up), to come up llek-üm to raise, to cultivate, to grow (plants), to bring out S. G.
33. lliw $\sim$ llüw $\sim$ lluw IV to get merged, to get melted, to get liquefied lliw-üm to merge, to melt, to liquefy G.
34. llum IV to get hidden, to conceal oneself llum-üm to hide, to conceal K.
35. lüf Aj burned, roasted p-üm to burn, to set fire, to roast S. G.
36. lüg Aj white lüg-üm to whiten, to bleach, to lighten (in color) S.
37. lulu N noise, crash, upheaval lulu-m to make noise, to cause crashing K.
38. nag Av down, underneath nak-üm to go down, to lower, to descend K. S. G.
39. nel Aj loose, untied, unleashed nel-üm $\sim$ nelk-üm to release, to let go, to free S.
40. neng IV to move (oneself) neng-üm to move, to shake, to shudder, to direct K.
41. neyküe IV to unleash, to get free neykü-m to release, to let go, to free S.
@@
42. ngül IV to get gathered, to come together, to join ngül-üm to gather, to bring together, to collect, to harvest K. S.
@@
43. nong IV to get broken down, to get damaged, to get scorched nong-üm to spoil, to damage, to scorch K.
44. nor Aj straight, correct, right nor-üm to correct, to rectify, to straighten S.
45. ñam IV to get lost, to get astray, to get loose ñam-üm to lose, to release, to let go, to make disappear S. G.
46. ñif IV to dry out ñip-üm to dry S.
47. ñom Aj meek, gentle, tame, docile, obedient ñom-üm to tame, to domesticate, to tame K. S.
48. pelof N light, clarity pelo-m to illuminate, to light up, to turn on the light S.
@@
49. pewü IV to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m to measure, to reach out using the arms, to twist, to turn K.
50. piwü IV to dry out piwü-m to dry (up) K. S.
51. pod N dirt, filth, uncleanliness pod-üm to dirty, to soil S.
52. poñü N potato poñü-m feed with potatoes K.
@@
53. püd Aj thick, dense, scattered, dispersed püd-üm to spread, to scatter, to thicken, to extend, to spread out, to stretch, to ration, to distribute K. S.
54. pümü Aj tense, tight, taut, strained pümü-m to stretch, to tighten, to tense K.
55. püna IV to get stuck, to get glued, to get adhered püna-m to stick, to glue, to adhere S.
@@
56. püra IV to ascend, to go up, to climb püra-mg to lift, to raise, to pick up, to keep in mind, to remember, to make evident, to highlight K. S. G.
57. püsha IV to run out, to be finished püsha-m to finish, to end, to spend, to use up, to exhaust G.
58. pütre Aj burning, red-hot pütre-m to smoke, to fumigateh K.
@@
59. puw IV to arrive, to reach, to achieve, to stay, to remain puw-üm to send, to deliver, to convey, to serve K.
60. rüngü IV to be pulverised, to be ground, to be crushed rüngü-m to grind, to mill, to crush S.
@@
61. traf Aj together, joined, gathered, fitted trap-üm to gather, to collect, to put together, to grab, to catch, to pick up K. S.
@@
62. tral IV to get hit, to get bumped, to get dented, to get bruised, to collide, to crash, to clash tral-üm to hit, to strike, to beat K.
@@
63. tran Aj fallen (especially referring to trees) tran-üm to make fall, to knock down, to knock over, to topple, to lay down G.
64. trar IV to ooze, to suppurate, to secrete trar-üm to drain the pus S.
65. trem Aj grown, adult, mature trem-üm to raise, to breed K. S. G.
66. trof IV to burst, to break, to crack, to snap trop-üm to split, to break S.
67. trong Aj thick, dense trong-üm to roof (houses)i K.
@@
68. trür Aj even, level, tight, composed trür-üm to equalise, to compensate, to balance, to indemnify K. S.
69. tüng IV to calm down, to take time tüng-üm to leave alone, not to disturb K.
70. üng Av while üng-üm to wait K.
71. üre $\sim$ ürü IV to get wet üre-m $\sim$ ürü-m to wet, to moisten, to soak K. S.
72. ürkü IV to get tired ürkü-m to tire, to tire out G.
73. üy $\sim$ üyü IV to catch fire, to ignite üy-üm $\sim$ üyü-m to set fire, to ignite K. S.
74. üy N name üy-üm to name, to denominate S.
75. wadkü IV to boil wadkü-m to boil, to bring to a boil K. S. G.
76. wallü IV to get surrounded, to get encompassed, to get completed wallü-m to tell or recount in detail K.
77. wef IV to appear, to show itself wep-üm to make appear, to find, to show G.
78. welli IV to get empty welli-m to empty, to vacate K. S.
79. wim Aj accustomed, used to, meek, tame, domesticated wim-üm to make accustomed, to domesticate, to tame S.
80. wüda IV to divide, to split up wüda-m to divide, to distribute, to share, to split K. S.
81. wünü IV to get stretched, to get elongated wünü-m to stretch, to lengthen S.
82. yall N son (of a man) yall-üm to multiply, to engender S.
83. yewe IV to feel embarrassed, to be ashamed yewe-m to embarrass G.
84. yung Aj pointed, sharp yung-üm to sharpen, to make pointed K.
a The “Collected from” column indicates which source contains examples of the root with the causative suffix.
K. = Wilhelm et al., 1930 (autobiography of longko Paskwal Koña); S. = Smeets, 2008 (A Grammar of Mapuche); G. = Golluscio, 2007 (Gramática descriptiva del mapuche).
@@
b The Mapuche use the term chaw-üm ‘fathering’, or more literally ‘to cause to father’, to refer to ‘incubating eggs’.